# Pulse Profiles of Accreting Neutron Stars from GRMHD Simulations

### v1

@@ Abstract
The pulsed X-ray emission from the neutron star surface acts as a window to study the state of matter in the neutron star interior. For accreting millisecond pulsars, the surface X-ray emission is generated from the ‘hotspots’, which are formed as a result of magnetically channeled accretion flow hitting the stellar surface. The emission from these hotspots is modulated by stellar rotation giving rise to pulsations. Using global three-dimensional general relativistic magnetohydrodynamic (GRMHD) simulations of the star-disk system, we investigate the accretion hotspots and the corresponding X-ray pulse properties of accreting millisecond pulsars with dipolar magnetic fields. The accretion spot morphologies in our simulations are entirely determined by the accretion columns and vary as a function of the stellar magnetic inclination. For lower inclinations, the hotspots are shaped like crescents around the magnetic axis. As we increase the inclination angle, the crescents transform into elongated bars close to the magnetic pole. We model the X-ray pulses resulting from the accretion hotspots using general-relativistic ray tracing calculations and quantify the root mean square variability of the pulsed signal. The pulse amplitudes obtained from our simulations usually range between $1-12\%$ rms and are consistent with the values observed in accreting millisecond pulsars. We find that the turbulent accretion flow in the GRMHD simulations introduces significant broadband variability on a timescale similar to the stellar rotational period. We also explore the impact of electron scattering absorption and show that, along with being a key factor in determining the pulse characteristics, this also introduces significant additional variability and higher harmonics in the bolometric light curve of the accreting sources.
@@ 1 Introduction
PPM can also be applied to AMXPs (Salmi et al., 2018; Dorsman et al., 2024), which is an important science goal for future large-area X-ray spectral timing telescopes (Watts et al., 2016), but this brings new challenges. The surface emission mechanism and hotspot properties are different to the RMPs, being formed by the channeled accretion flow rather than return current heating. This indicates that the surface pattern models used in RMP PPM need to be modified for the AMXPs. And unlike the RMPs, AMXP pulse profiles are variable, with pulse fractional amplitudes, harmonic content, and pulse phases varying significantly over time (see e.g. Hartman et al., 2008; Patruno et al., 2009b, 2010).
@@ 2.1 GRMHD Simulations
We use the total variation diminishing Lax-Friedrich scheme for fluxes along with a Piecewise Parabolic reconstruction scheme and second-order modified Euler time-stepper. Our simulation domain extends from $-400\,r_{g}$ to $400\,r_{g}$ in $x$, $y$, and $z$ [^1] For the run with $\chi_{\rm star}=30^{\circ}$, we extend the domain along the z-axis such that $z$ $\in[-800,800]\,r_{g}$. Here $\chi_{\rm star}$ represents the angle between the magnetic axis and the rotation axis ($\Omega_{\rm star}$).. The base resolution of the domain is $128\times 128\times 128$ cells with 7 levels of AMR (adaptive mesh refinement). This allows us to resolve dynamically interesting regions with an effective resolution of $8192^{3}$ cells. Unless mentioned otherwise, we always use geometric units with $\textrm{G}=c=1$.
The stellar radius, magnetic field strength (denoted by the dipole moment $\mu$), and rotational frequency for all the GRMHD runs are set to $R_{\star}=4\,r_{g}$, $\mu=20$, and $\Omega_{\rm star}=0.03$ respectively. We extract all the relevant quantities on a spherical shell with radius $r=4.3\,r_{g}$. This ensures that the interpolation stencil stays clear of the stellar interior where the solution is excised. Thus, the equatorial radius of the neutron star ($R_{\rm eq}$) is always assumed to be $4.3\,r_{g}$ for the ray-tracing calculations. In order to extract observable signatures from the GRMHD simulations, we scale our simulations to physical units following Das et al. (2022). Assuming $R_{\star}=10$ km, allow us to convert the stellar parameters to the following physical units, $M=1.69\,{M}_{\odot}$, $R_{\rm eq}$ = 10.75 km, which sets $\Omega_{\mathrm{star}}=572.56$ Hz. The pulses are computed for a range of observer frequencies, $\nu_{\rm{obs}}\in[2.42\times 10^{15},1.23\times 10^{19}]$ Hz ([0.01, 50.87] KeV).
@@ 2.2.1 RAPTOR
Here, $\lambda$ is the affine parameter; $x^{\alpha}$, $k^{\alpha}$, and $\Gamma_{\mu v}^{\alpha}$ are the position, contravariant wave vector, and Christoffel symbols, respectively. The initial wave vectors are constructed following Cunningham & Bardeen (1972). The initial positions of the photon are always set by the location of the virtual camera which is set to ($r,\theta,\phi$) = $(10^{4}\,r_{g},i_{\rm obs},0)$, where $i_{\rm obs}$ represents the observer inclination. The numerical integration is performed using a 4th-order Runge-Kutta method. In this ray-tracing method, the photons are traced backward from the virtual camera, and thus, also known as the backward integration method. In RAPTOR, we assume a fast-light approximation where the geodesic’s time coordinate is ignored. The impact of ignoring time delays on the pulse profiles is shown for the test cases in Section 2.2.3. The base resolution of the virtual camera is set to $400\times 400$ pixels with 2 levels of AMR resolving the hotspots with an effective of $800\times 800$ pixels [^2] We have performed a convergence study which showed that the difference in spectrum stays within $0.25\%$ for the runs with $1600\times 1600$ pixels and $400\times 400$ pixels with 2 AMR levels. (See Davelaar & Haiman 2022 for a detailed description of the AMR implementation in RAPTOR).
@@ 2.2.2 X-PSI
(Bogdanov et al., 2019) where the primed quantities are measured in the local surface comoving frame (so at the stellar surface $I^{\prime}_{\nu^{\prime}}$ is equivalent to $I_{\nu}$ mentioned in Equation (6)), $D$ is the distance to the star, $\sigma^{\prime}$ is the emission angle relative to the surface normal, $dS^{\prime}$ is a surface area element, and $\delta$ is the Doppler factor that depends on the NS surface velocity and the angle between the velocity and the direction of the emitted photon. The effect of time delays due to different photon paths is usually accounted for in the X-PSI calculation (unlike in RAPTOR), but in some cases in Section 2.2.3 it is turned off to match RAPTOR results. [^3] Ignoring the geodesic’s time coordinate in RAPTOR affects also the observed number of photons per second even for a single photon path. Thus, when ignoring time delays, we also neglect one of the Doppler factors in Equation (11), coming from the photon arrival time contraction (see e.g., Rybicki & Lightman, 1979; Poutanen & Beloborodov, 2006). The calculations were based on X-PSI version v1.2.1444https://github.com/xpsi-group/xpsi modified additionally to allow surface temperature maps from the GRMHD simulations.
@@ 3.1 GRMHD Simulations
$F_{\rm m}$ $={T^{\text{MA}}}^{r}_{t}=(\rho_{0}+u_{g}+p_{g})u^{r}u^{t}\,.$ (13)
Here, $F_{\rm m}$ is the total matter-energy flux; ${T^{\text{MA}}}^{r}_{t}$ is the matter component of the stress-energy tensor (McKinney et al., 2012); $T_{\rm eff}$ and $\sigma_{\rm SB}$ are the effective black-body temperature and Stefan-Boltzmann constant respectively. Finally, assuming $\dot{M}=1$ [code units] corresponds to $1\%\dot{M}_{\rm Edd}$ (radiative efficiency = 0.1), we compute the temperature in physical (cgs) units. Here, $\dot{M}_{\rm Edd}$ represents the mass accretion rate corresponding to the Eddington luminosity.
@@
Figure 3: Variability of temperature profiles and spot shapes for $\chi_{\mathrm{star}}=60^{\circ}$ around $1200r_{\textrm{g}}/c$ ($\approx 10$ ms) apart. The variation in the high-temperature regions ($T_{\mathrm{eff}}>2.5\times 10^{7}$ K) is outlined with the green contours. A movie showing the variability is available on Zenodo at https://doi.org/10.5281/zenodo.10863742 (Das et al., 2024).
@@ 3.2.2 Variability
Figure 7: rms normalized power spectrum of (a) Luminosity ($L$) and (b) $\dot{M}$ for $\chi_{\mathrm{star}}=30^{\circ}$ (blue), $60^{\circ}$ (orange), and $90^{\circ}$ (green). The black and grey dashed lines show $\nu_{0}=\Omega_{\mathrm{star}}$ and $\nu_{1}=2\Omega_{\mathrm{star}}$ respectively.
@@ 3.3 Accretion Column
Figure 9 illustrates the variation in pulse shape as a function of mass accretion rate for $\chi_{\rm star}=60^{\circ}$ and $i_{\rm obs}$$=50^{\circ}$. We begin with a mass accretion rate of $0.01\%\dot{M}_{\rm Edd}$ and gradually increase it to study the role of electron scattering absorption as the system transitions from quiescence to a moderate accretion rate. For a given GRMHD simulation, scaling the surface $\dot{M}$ to higher physical accretion rates is accompanied by an increased stellar magnetic field strength as given by Equation 6 in Das et al. (2022). This ensures a constant magnetospheric radius regardless of the physical mass accretion rate scaling. As we increase the mass accretion rate, two effects come into play: on the one hand, the surface-emission increases; on the other hand, the optical depth also rises, which impacts the pulse profiles.
@@ 4 Discussion &\&& Conclusions
Although complex spot shapes involving, e.g., crescents have been applied in pulse profile modeling studies of rotation-powered pulsars (Riley et al., 2019), to date, PPM studies for accretion powered pulsars have been adopting circular hotspot geometries (see e.g. Salmi et al., 2018, 2021; Ahlberg et al., 2024; Dorsman et al., 2024). The family of crescent and elongated band shaped hotspots resulting from the global GRMHD simulations, along with the impact these complex spot geometries have on the simulated pulse profiles, strongly suggests the need to move beyond the assumption of circular spots in case of accretion-powered pulsars.
@@
Our current study treats scattering only as (monochromatic) absorption opacity which dims the signal from the pulse but ignores line-of-sight scattering and energy scattering. We have therefore focused on results regarding the bolometric pulse properties. We note, however, that our predicted pulse amplitudes can be slightly overestimated, since we treat the initial hotspot spectrum and beaming as isotropic blackbody here (see e.g. Figure 4 of Poutanen & Gierliński 2003 for the effect). More accurate modeling of the neutron star atmosphere, subsequent scattering, as well as incorporating the background emission from the disk are necessary to make predictions on the pulse properties with energy. In particular, determining the disk emission ‘background’ is an important task for mass-radius inference of such sources (Patruno et al., 2009a; Kini et al., 2024b; Dorsman et al., 2024). This requires more advanced radiation transfer calculations which we leave for future work. Furthermore, it will be very interesting to sample pulse profiles for a number of simulations with varying magnetospheric radii. This will allow us to model changes in pulse profiles with varying mass accretion rate onto a given source (e.g. Hartman et al., 2008; Patruno et al., 2009b, 2010).
@@ Appendix A Accretion Columns
Figure 11 shows the volume rendering of density for $\chi_{\rm star}=60^{\circ}$ (top panel) and $\chi_{\rm star}=90^{\circ}$ (bottom panel). The solid lines show the magnetic field, and the color gradient in the lines shows the magnitude of the field.
@@
Figure 11: Volume rendering of density for $\chi_{\rm star}=60^{\circ}$ (top panel) and $\chi_{\rm star}=90^{\circ}$ (bottom panel). The solid lines show the corresponding stellar magnetic fieldlines for each case.
Both panels clearly show the formation of symmetric accretion columns across the magnetic equator (magnetic axis in the $z-x$ plane, axes shown in the lower left corner of the plot) close to the $\mu-y$ plane. The sphere at the center of the plots shows the density map at the stellar surface.
@@ Appendix B Power Spectrum Densities of the bolometric light curve with absorption
Figure 12: rms normalized PSDs of the bolometric luminosity with electron scattering absorption for different observer inclinations, $i_{\rm obs}=50^{\circ}$ (left) and $i_{\rm obs}=70^{\circ}$ (right) for $\chi_{\rm star}=60^{\circ}$ at $1\%\dot{M}_{\rm Edd}$. The blue and orange curves shows the PSDs resulting from the surface hotspots alone and surface emission with electron scattering absorption respectively. The black and grey dashed lines with reducing opacities show $\nu_{0}=\Omega_{\mathrm{star}}$, $\nu_{1}=2\Omega_{\mathrm{star}}$, $\nu_{2}=3\Omega_{\mathrm{star}}$, $\nu_{3}=4\Omega_{\mathrm{star}}$ and $\nu_{4}=5\Omega_{\mathrm{star}}$ respectively.

### v2

@@ Abstract
The pulsed X-ray emission from the neutron star surface acts as a window to study the state of matter in the neutron star interior. For accreting millisecond pulsars, the surface X-ray emission is generated from the ‘hotspots’ formed due to the magnetically channeled accretion flow hitting the stellar surface. The emission from these hotspots is modulated by stellar rotation giving rise to pulsations. Using global three-dimensional general relativistic magnetohydrodynamic (GRMHD) simulations of the star-disk system, we investigate the accretion hotspots and the corresponding X-ray pulse properties of accreting millisecond pulsars with dipolar magnetic fields. The accretion spot morphologies in our simulations are entirely determined by the accretion columns and vary as a function of the stellar magnetic inclination. For lower magnetic inclinations, the hotspots are shaped like crescents around the magnetic axis and are transformed into elongated bars for higher inclinations. We model the X-ray pulses resulting from the simulated hotspots using general-relativistic ray tracing calculations and quantify the variability of the pulsed signal. The pulse amplitudes in our simulations usually range between $1-12\%$ rms and are consistent with the observed values. We find that the turbulent accretion flow in the GRMHD simulations introduces significant broadband variability on a timescale similar to the stellar rotational period. We also explore the impact of electron scattering absorption and show that along with being a key factor in determining the pulse characteristics, this also introduces significant additional variability and higher harmonics in the bolometric light curve of the accreting sources.
@@ 1 Introduction
PPM can also be applied to AMXPs (Salmi et al., 2018; Dorsman et al., 2025), which is an important science goal for future large-area X-ray spectral timing telescopes (Watts et al., 2016), but this brings new challenges. The surface emission mechanism and hotspot properties are different to the RMPs, being formed by the channeled accretion flow rather than return current heating. This indicates that the surface pattern models used in RMP PPM need to be modified for the AMXPs. And unlike the RMPs, AMXP pulse profiles are variable, with pulse fractional amplitudes, harmonic content, and pulse phases varying significantly over time (see e.g. Hartman et al., 2008; Patruno et al., 2009b, 2010).
@@ 2.1 GRMHD Simulations
Unless mentioned otherwise, we always use geometric units with $\textrm{G}=c=1$. The stellar radius, magnetic field strength (denoted by the dipole moment $\mu$), and rotational frequency for all the GRMHD runs are set to $R_{\star}=4\,r_{g}$, $\mu=20\,\mu_{cgs}$, and $\Omega_{\rm star}=0.03\,c/r_{g}$ respectively, where $\mu_{cgs}$ is $\sqrt{4\pi\rho_{\rm cgs}c^{2}}\,r_{g}^{3}$. Finally, assuming $\dot{M}=1$ [code units] corresponding to $1\%\dot{M}_{\rm Edd}$ results in a physical density scaling of $\rho_{\rm cgs}=1.27\times 10^{-5}$ g cm-3 for a radiative efficiency of 0.1. Here, $\dot{M}_{\rm Edd}$ denotes the mass accretion rate corresponding to the Eddington luminosity. We extract all the relevant quantities on a spherical shell with radius $r=4.3\,r_{g}$. This ensures that the interpolation stencil stays clear of the stellar interior where the solution is excised. Thus, the equatorial radius of the neutron star ($R_{\rm eq}$) is always assumed to be $4.3\,r_{g}$ for the ray-tracing calculations. In order to extract observable signatures from the GRMHD simulations, we scale our simulations to physical units following Das et al. (2022). Assuming $R_{\star}=10$ km, allow us to convert the stellar parameters to the following physical units, $M=1.69\,{M}_{\odot}$, $R_{\rm eq}$ = 10.75 km, which sets $\Omega_{\mathrm{star}}=572.56$ Hz. The pulses are computed for a range of observer frequencies, $\nu_{\rm{obs}}\in[2.42\times 10^{15},1.23\times 10^{19}]$ Hz ([0.01, 50.87] KeV).
@@ 2.2.1 RAPTOR
Here, $\lambda$ is the affine parameter; $x^{\alpha}$, $k^{\alpha}$, and $\Gamma_{\mu v}^{\alpha}$ are the position, contravariant wave vector, and Christoffel symbols, respectively. The initial wave vectors are constructed following Cunningham & Bardeen (1972). The initial positions of the photon are always set by the location of the virtual camera which is set to ($r,\theta,\phi$) = $(10^{4}\,r_{g},i_{\rm obs},0)$, where $i_{\rm obs}$ represents the observer inclination. The numerical integration is performed using a 4th-order Runge-Kutta method. In this ray-tracing method, the photons are traced backward from the virtual camera, and thus, also known as the backward integration method. In RAPTOR, we assume a fast-light approximation where the geodesic’s time coordinate is ignored. The impact of ignoring time delays on the pulse profiles is shown for the test cases in Section 2.2.3. The base resolution of the virtual camera is set to $400\times 400$ pixels with 2 levels of refinement resolving the hotspots with an effective of $800\times 800$ pixels [^1] We have performed a convergence study which showed that the difference in spectrum stays within $0.25\%$ for the runs with $1600\times 1600$ pixels and $400\times 400$ pixels with 2 refinement levels. (See Davelaar & Haiman 2022 for a detailed description of the camera refinement implementation in RAPTOR). The camera in RAPTOR is a block-based quadtree, for each block, we first compute the emission at the lowest refinement level. Then by taking the gradient of the specific intensity sharp features are identified. If a sharp feature is present, the block will split into four, with two new blocks in each direction, each containing the same number of pixels. This is done iteratively until the sharp features are either fully resolved or when a user-defined maximum depth is achieved. This way is reminiscent of the adaptive mesh refinement strategies exploited in the GRMHD codes.
@@ 2.2.2 X-PSI
(Bogdanov et al., 2019) where the primed quantities are measured in the local surface comoving frame (so at the stellar surface $I^{\prime}_{\nu^{\prime}}$ is equivalent to $I_{\nu}$ mentioned in Equation (6)), $D$ is the distance to the star, $\sigma^{\prime}$ is the emission angle relative to the surface normal, $dS^{\prime}$ is a surface area element, and $\delta$ is the Doppler factor that depends on the NS surface velocity and the angle between the velocity and the direction of the emitted photon. The effect of time delays due to different photon paths is usually accounted for in the X-PSI calculation (unlike in RAPTOR), but in some cases in Section 2.2.3 it is turned off to match RAPTOR results. [^2] Ignoring the geodesic’s time coordinate in RAPTOR affects also the observed number of photons per second even for a single photon path. Thus, when ignoring time delays, we also neglect one of the Doppler factors in Equation (11), coming from the photon arrival time contraction (see e.g., Rybicki & Lightman, 1979; Poutanen & Beloborodov, 2006). The calculations were based on X-PSI version v1.2.1333https://github.com/xpsi-group/xpsi modified additionally to allow surface temperature maps from the GRMHD simulations.
@@ 3.1 GRMHD Simulations
$F_{\rm m}$ $={T^{\text{MA}}}^{r}_{t}=(\rho_{0}+u_{g}+p_{g})u^{r}u_{t}\,.$ (13)
Here, $F_{\rm m}$ is the total matter-energy flux; ${T^{\text{MA}}}^{r}_{t}$ is the matter component of the stress-energy tensor (McKinney et al., 2012); $T_{\rm eff}$ and $\sigma_{\rm SB}$ are the effective black-body temperature and Stefan-Boltzmann constant respectively. Finally, assuming $\dot{M}=1$ [code units] corresponds to $1\%\dot{M}_{\rm Edd}$ (radiative efficiency = 0.1), we compute the temperature in physical (cgs) units.
@@
Figure 3: Variability of temperature profiles and spot shapes for $\chi_{\mathrm{star}}=60^{\circ}$ around $1200r_{\textrm{g}}/c$ ($\approx 10$ ms) apart. The variation in the high-temperature regions ($T_{\mathrm{eff}}>2.5\times 10^{7}$ K) is outlined with the green contours. An animated version of the figure, illustrating the temperature evolution over $\rm 83\,ms$ with a cadence of $\rm 0.17\,ms$, is available at: https://youtu.be/7QD3z1SKESE.
@@ 3.2.2 Variability
Figure 7: RMS-normalized power spectrum of (a) Luminosity ($L$) and (b) $\dot{M}$ for $\chi_{\mathrm{star}}=30^{\circ}$ (blue), $60^{\circ}$ (orange), and $90^{\circ}$ (green). The black and grey dashed lines show $\nu_{0}=\Omega_{\mathrm{star}}$ and $\nu_{1}=2\Omega_{\mathrm{star}}$ respectively.
@@ 3.3 Accretion Column
Figure 9 illustrates the variation in pulse shape as a function of mass accretion rate for $\chi_{\rm star}=60^{\circ}$ and $i_{\rm obs}$$=50^{\circ}$. We begin with a mass accretion rate of $0.01\%\dot{M}_{\rm Edd}$ and gradually increase it to study the role of electron scattering absorption as the system transitions from quiescence to a moderate accretion rate. For a given GRMHD simulation, scaling the surface $\dot{M}$ to higher physical accretion rates is accompanied by an increased stellar magnetic field strength as given by Equation 6 in Das et al. (2022). Following the unit conversion described in section 2, a mass accretion rate ranging from $0.01\%\dot{M}_{\rm Edd}$ to $1\%\dot{M}_{\rm Edd}$ corresponds to a magnetic field strength ranging from $\rm 3.78\times 10^{7}\,G$ to $\rm 3.78\times 10^{8}\,G$. This ensures a constant magnetospheric radius regardless of the physical mass accretion rate scaling. As we increase the mass accretion rate, two effects come into play: on the one hand, the surface emission increases; on the other hand, the optical depth also rises, which impacts the pulse profiles.
@@ 4 Discussion &\&& Conclusions
Although complex spot shapes involving, e.g., crescents have been applied in pulse profile modeling studies of rotation-powered pulsars (Riley et al., 2019), to date, PPM studies for accretion powered pulsars have been adopting circular hotspot geometries (see e.g. Salmi et al., 2018, 2021; Ahlberg et al., 2024; Dorsman et al., 2025). The family of crescent and elongated band shaped hotspots resulting from the global GRMHD simulations, along with the impact these complex spot geometries have on the simulated pulse profiles, strongly suggests the need to move beyond the assumption of circular spots in case of accretion-powered pulsars.
@@
Our current study treats scattering only as (monochromatic) absorption opacity which dims the signal from the pulse but ignores line-of-sight scattering and energy scattering. We have therefore focused on results regarding the bolometric pulse properties. Including scattering into the line of sight would predominantly affect the southern hotspot which is strongly absorbed by the accretion disk matter starting at Eddington fractions of $1\%$. Due to directional scattering, we thus expect the dips of the first harmonic to be less pronounced and broadened compared to the case studied here. Furthermore, we note that our predicted pulse amplitudes can be slightly overestimated since we treat the initial hotspot spectrum and beaming as isotropic blackbody here (see e.g. Figure 4 of Poutanen & Gierliński 2003 for the effect). More accurate modeling of the neutron star atmosphere, subsequent scattering, as well as incorporating the background emission from the disk are necessary to make predictions on the pulse properties with energy and assess the effect of line-of-sight scattering. In particular, determining the disk emission ‘background’ is an important task for mass-radius inference of such sources (Patruno et al., 2009a; Kini et al., 2024b; Dorsman et al., 2025). This requires more advanced radiation transfer calculations which we leave for future work. Furthermore, it will be very interesting to sample pulse profiles for a number of simulations with varying magnetospheric radii. This will allow us to model changes in pulse profiles with varying mass accretion rate onto a given source (e.g. Hartman et al., 2008; Patruno et al., 2009b, 2010).
@@ Appendix A Accretion Columns
Figure 11 shows the volume rendering of density for $\chi_{\rm star}=60^{\circ}$ (top panel) and $\chi_{\rm star}=90^{\circ}$ (bottom panel). The solid lines show the magnetic field, and the color gradient in the lines shows the magnitude of the field. Both panels clearly show the formation of symmetric accretion columns across the magnetic equator (magnetic axis in the $z-x$ plane, axes shown in the lower left corner of the plot) close to the $\mu-y$ plane. The sphere at the center of the plots shows the density map at the stellar surface.
@@
Figure 11: Volume rendering of density for $\chi_{\rm star}=60^{\circ}$ (left panel) and $\chi_{\rm star}=90^{\circ}$ (right panel). The solid lines show the corresponding stellar magnetic fieldlines for each case.
@@ Appendix B Power Spectrum Densities of the bolometric light curve with absorption
Figure 12: RMS-normalized PSDs of the bolometric luminosity with electron scattering absorption for different observer inclinations, $i_{\rm obs}=50^{\circ}$ (left) and $i_{\rm obs}=70^{\circ}$ (right) for $\chi_{\rm star}=60^{\circ}$ at $1\%\dot{M}_{\rm Edd}$. The blue and orange curves represent the PSDs resulting from the surface hotspots alone and surface emission with electron scattering absorption respectively. The black and grey dashed lines, with decreasing opacity, indicate $\nu_{0}=\Omega_{\mathrm{star}}$, $\nu_{1}=2\Omega_{\mathrm{star}}$, $\nu_{2}=3\Omega_{\mathrm{star}}$, $\nu_{3}=4\Omega_{\mathrm{star}}$ and $\nu_{4}=5\Omega_{\mathrm{star}}$ respectively.